OpenStudy (anonymous):

Let a,b,c be integers such that the average of the numbers 7a−9, 7b−9, 7c−9is2014.Prove that the average of the numbers a,b,c is a perfect square

ganeshie8 (ganeshie8):

kids problem :P

OpenStudy (anonymous):

:P well assume you dnt know the avg solve it

OpenStudy (anonymous):

lets say solve it theoritical

ganeshie8 (ganeshie8):

if i dnt even knw avg, how would u expect me knw theorotical lol

OpenStudy (anonymous):

humm number theory :O

OpenStudy (anonymous):

im trying lol , found it interested :3

ganeshie8 (ganeshie8):

7a−9 + 7b−9 + 7c−9 = 3*2014

ganeshie8 (ganeshie8):

is that the starting step ?

OpenStudy (anonymous):

idk xD

OpenStudy (anonymous):

so it means a+b+c/3 should be of the form n^2

ganeshie8 (ganeshie8):

yep

OpenStudy (anonymous):

nvm lol , i wake up suddly thinking of this xD ill go back to slp

ganeshie8 (ganeshie8):

:) i thought u had some nice way to solve this... without using average thing... .

OpenStudy (anonymous):

unless 7(a+b+c)-9 /3 = 2014 its weird cuz 2014=7(..)+5 :-\

ganeshie8 (ganeshie8):

you get (7(a+b+c)-27)/3 = 2014 right ?

OpenStudy (anonymous):

O.O yeah

ganeshie8 (ganeshie8):

(7(a+b+c)-27)/3 = 2014 7(a+b+c)/3 - 9 = 2014 7(a+b+c)/3 = 2023 (a+b+c)/3 = 289 = 17^2 QED, go back to slp :D

OpenStudy (anonymous):

haha ok good night

OpenStudy (anonymous):

btw im not sure if i could sleep peacefully until i figure out something else :-\ but its better to slp :3